# Enhancing Cognition and Explainability of Multimodal Foundation Models with Self-Synthesized Data

### v1

@@ Abstract
Large multimodal models (LMMs) have shown impressive capabilities in a wide range of visual tasks. However, they often struggle with fine-grained visual reasoning, failing to identify domain-specific objectives and provide justifiable explanations for their predictions. To address this, we propose a novel visual rejection sampling framework to improve the cognition and explainability of LMMs using self-synthesized data. Specifically, visual fine-tuning requires images, queries, and target answers. Our approach begins by synthesizing interpretable answers that include human-verifiable visual features. These features are based on expert-defined concepts, carefully selected based on their alignment with the image content. After each round of fine-tuning, we apply a reward model-free filtering mechanism to select the highest-quality interpretable answers for the next round of tuning. This iterative process of data synthesis and fine-tuning progressively improves the model’s ability to generate accurate and reasonable explanations. Experimental results demonstrate the effectiveness of our method in improving both the accuracy and explainability of specialized visual classification tasks.
@@ 1 Introduction
Figure 2: Examples of synthesized answers for query Q. Training with the first two types leads to shortcut learning or overgeneralization.
@@
In summary, our contributions are threefold: (1) We propose a novel framework that improves LMMs’ interpretable visual classification abilities without requiring extensive manual labeling, (2) We introduce an information-theoretic approach to select interpretable visual concepts for each image and a reward model-free filtering method to ensure high-quality data selection from synthesized outputs, and (3) We develop an iterative process of data synthesis and model fine-tuning to progressively enhance LMMs’ cognitive abilities and explainability.
@@ 3.1 Overview
Our approach leverages synthesized data for visual fine-tuning to enhance both the cognitive ability and explainability of LMMs, as illustrated in Figure 3. There are two major steps:
1. Image-level Visual Concept Selection: Given an image and its label, our first task is to extract a set of image-specific concepts that explain the connection between the image and label. We propose a selection method that identifies a subset of concepts most relevant to the image content while ensuring the subset is concise. Using these concepts, we prompt the LMM to rewrite them into textual answers that explains the label. These answers are used for the first round of fine-tuning.
2. Reward Model-Free Rejection Sampling: After the initial fine-tuning, the model can generate synthetic answers, which can be used for next-round fine-tuning, but their quality still varies. To filter out low-quality explanations and only select the best quality one as training data, we use previously selected concepts as filtering criteria, quantifying alignment between explanations and the concepts. The synthetic answer with the best-aligned explanation is then selected and combined with the corresponding image and query, forms a new data pair for subsequent rounds of fine-tuning.
@@ 3.3 Step 2: Reward Model-Free Rejection Sampling
After the initial round of fine-tuning with explainable visual query-answer pairs, the LMMs has significantly improved their ability to generate reasonable explanations. This improvement allows us to leverage the current fine-tuned model to generate new data for subsequent training rounds. However, the quality of newly generated data can vary considerably in terms of label accuracy and explanation quality. Training on low-quality data could lead to performance degradation. To address this issue, we propose a rejection sampling technique that filters out low-quality outputs.
@@
In summary, our framework, outlined in Algorithm 1, enhances the model’s fine-grained classification capabilities by iteratively fine-tuning on diverse, high-quality, synthesized visual classification query-answer pairs. Step 1 identifies the most informative concepts in the image, while Step 2 selects the explanations that best align with these concepts. This two-step approach improves both the accuracy and interpretability of the model’s predictions, enabling it to perform more effectively in complex visual classification tasks.
@@
1:Dataset $\mathcal{D}=\{(X_{i},c_{i})\}_{i=1}^{N}$, Concept Sets $\mathcal{Z}=\{Z_{i}\}_{i=1}^{|C|}$ of every label, Pre-trained LMM $f_{\theta}^{0}$, Synthesized query set $Q=\{q_{i}\}_{i=1}^{N}$, Descriptive query $q_{d}$
@@ 4.1 Experiments Settings
4.2 RQ1: Training on our synthesized data improves classification
@@
Our proposed method achieves higher accuracy than both NL and L+GE baselines, as shown in Table 1. While the baselines, particularly NL, exhibit overfitting, our method improves accuracy with each iteration, demonstrating its effectiveness in fine-grained classification tasks. Our approach’s resilience to overfitting stems from the use of rejection sampling, which generates more diverse training dataset. This expanded data pool enhances the model’s generalization capabilities. Moreover, training with our dataset, which includes specific visual features, helps the model learn more detailed visual knowledge, contributing to further gains in classification performance. In contrast, the L+GE baseline struggles with certain datasets, notably HAM10000, likely due to irrelevant information in general explanations impeding effective learning. Moreover, we assessed the general ability of the models using the MMMU metric (Yue et al., 2024) evaluated by the LMMs-eval tool (Zhang et al., 2024). The results indicate that all trained models maintain comparable general abilities to the base model, with negligible degradation. Note that the LMMs-eval tool does not support the LLaVA-Med model; therefore, the general ability metric is not reported. More experiment results can be found in the appendix.
@@ 4.5 RQ4: Visualization of Visual Focus of LMMs
To interpret our model’s predictions, we visualize the outputs of LMMs using a gradient-based explanation method (Wu et al., 2023). This approach identifies the image regions most influential in generating the model’s answer by computing gradients of answer token probabilities with respect to image pixels. Figure 7 presents case examples demonstrating that training on our synthesized answers with cognitive explanations can effectively prevent the model from learning spurious features, whereas directly linking images to naive labels may result in shortcut learning.
@@ 5 Related Work: Self-improvement and Data Synthesis
In the field of LLMs, self-generated data has become a powerful tool for enhancing model performance (Liu et al., 2024d). Researchers have explored various techniques, such as rejection sampling (Touvron et al., 2023; Dubey et al., 2024), self-rewarding (Yuan et al., 2024), and self-play (Chen et al., 2024c), enabling models to improve using synthesized data. Recent studies (Snell et al., 2024; Brown et al., 2024) have proposed the “inference-time scaling law”, suggesting that increasing inference samples size boosts the likelihood of generating high-quality data. Data synthesis techniques have also been applied to LMMs, improving general vision tasks like visual question answering (Zhou et al., 2024a; Deng et al., 2024; Gao et al., 2023; Zhou et al., 2024b) and enhancing instruction-following capabilities (Luo et al., 2024). Our work extends this line of research by focusing on domain-specific visual classification to enable effective visual assistance in professional tasks. In contrast to existing methods, we address the unique challenges of data synthesis in specialized domains, thereby extending these techniques to support expert-driven applications.

### v2

@@ Abstract
Large Multimodal Models (LMMs), or Vision-Language Models (VLMs), have shown impressive capabilities in a wide range of visual tasks. However, they often struggle with fine-grained visual reasoning, failing to identify domain-specific objectives and provide justifiable explanations for their predictions. To address the above challenge, we propose a novel visual rejection sampling framework to improve the cognition and explainability of LMMs using self-synthesized data. Specifically, visual fine-tuning requires images, queries, and target answers. Our approach begins by synthesizing interpretable answers that include human-verifiable visual features. These features are based on expert-defined concepts, and carefully selected based on their alignment with the image content. After each round of fine-tuning, we apply a reward model-free filtering mechanism to select the highest-quality interpretable answers for the next round of tuning. This iterative process of synthetic data generation and fine-tuning progressively improves the model’s ability to generate accurate and reasonable explanations. Experimental results demonstrate the effectiveness of our method in improving both the accuracy and explainability of specialized visual classification tasks.
@@ 1 Introduction
Figure 2: Examples of synthetic answers for query Q. Training with the first two types leads to shortcut learning or overgeneralization.
@@
In summary, our contributions are threefold: (1) We propose a novel framework that improves LMMs’ interpretable visual classification abilities without requiring extensive manual labeling, (2) We introduce an information-theoretic approach to select interpretable visual concepts for each image and a reward model-free filtering method to ensure high-quality data selection from synthetic outputs, and (3) We develop an iterative process of data synthesis and model fine-tuning to progressively enhance LMMs’ cognitive abilities and explainability.
@@ 3.1 Overview
Our approach leverages synthetic data for visual fine-tuning to enhance both the cognitive ability and explainability of LMMs, as illustrated in Figure 3. There are two major steps:
1. Image-level Visual Concept Selection: Given an image and its label, our first task is to extract a set of image-specific concepts that explain the connection between the image and label. We propose a selection method that identifies a subset of concepts most relevant to the image content while ensuring the subset is concise. Using these concepts, we prompt the LMM to rewrite them into textual answers that explain the label. These answers are used for the first round of fine-tuning.
2. Reward Model-Free Rejection Sampling: After the initial fine-tuning, the model can generate synthetic answers, which can be used for next-round fine-tuning, but their quality still varies. To filter out low-quality explanations and only select the best quality one as training data, we use previously selected concepts as filtering criteria, quantifying alignment between explanations and the concepts. The synthetic answer with the best-aligned explanation is then selected and combined with the corresponding image and query, forming a new data pair for subsequent rounds of tuning.
@@ 3.3 Step 2: Reward Model-Free Rejection Sampling
After the initial round of fine-tuning with explainable visual query-answer pairs, the LMMs have significantly improved their ability to generate reasonable explanations. This improvement allows us to leverage the current fine-tuned model to generate new data for subsequent training rounds. However, the quality of newly generated data can vary considerably in terms of label accuracy and explanation quality. Training on low-quality data could lead to performance degradation. To address this issue, we propose a rejection sampling technique that filters out low-quality outputs.
@@
In summary, our framework, outlined in Algorithm 1, enhances the model’s fine-grained classification capabilities by iteratively fine-tuning on diverse, high-quality, synthetic visual classification query-answer pairs. Step 1 identifies the most informative concepts in the image, while Step 2 selects the explanations that best align with these concepts. This two-step approach improves both the accuracy and interpretability of the model’s predictions, enabling it to perform more effectively in complex visual classification tasks.
@@
1:Dataset $\mathcal{D}=\{(X_{i},c_{i})\}_{i=1}^{N}$, Concept Sets $\mathcal{Z}=\{Z_{i}\}_{i=1}^{|C|}$ of every label, Pre-trained LMM $f_{\theta}^{0}$, synthetic query set $Q=\{q_{i}\}_{i=1}^{N}$, Descriptive query $q_{d}$
@@ 4.1 Experiments Settings
4.2 RQ1: Training on our synthetic data improves classification
@@
Our proposed method achieves higher accuracy than both NL and L+GE baselines, as shown in Table 1. While the baselines, particularly NL, exhibit overfitting, our method improves accuracy with each iteration, demonstrating its effectiveness in fine-grained classification tasks. Our approach’s resilience to overfitting stems from the use of rejection sampling, which generates a more diverse training dataset. This expanded data pool enhances the model’s generalization capabilities. Moreover, training with our dataset, which includes specific visual features, helps the model learn more detailed visual knowledge, contributing to further gains in classification performance. In contrast, the L+GE baseline struggles with certain datasets, notably HAM10000, likely due to irrelevant information in general explanations impeding effective learning. Moreover, we assessed the general ability of the models using the MMMU metric (Yue et al., 2024) evaluated by the LMMs-eval tool (Zhang et al., 2024). The results indicate that all trained models maintain comparable general abilities to the base model, with negligible degradation. Note that the LMMs-eval tool does not support the LLaVA-Med model; therefore, the general ability metric is not reported. More experiment results can be found in the appendix.
@@ 4.5 RQ4: Visualization of Visual Focus of LMMs
To interpret our model’s predictions, we visualize the outputs of LMMs using a gradient-based explanation method (Wu et al., 2023). This approach identifies the image regions most influential in generating the model’s answer by computing gradients of answer token probabilities with respect to image pixels. Figure 7 presents case examples demonstrating that training on our synthetic answers with cognitive explanations can effectively prevent the model from learning spurious features, whereas directly linking images to naive labels may result in shortcut learning.
@@ 5 Related Work: Self-improvement and Data Synthesis
In the field of LLMs, self-generated data has become a powerful tool for enhancing model performance (Liu et al., 2024d). Researchers have explored various techniques, such as rejection sampling (Touvron et al., 2023; Dubey et al., 2024), self-rewarding (Yuan et al., 2024), and self-play (Chen et al., 2024c), enabling models to improve using synthetic data. Recent studies (Snell et al., 2024; Brown et al., 2024) have proposed the “inference-time scaling law”, suggesting that increasing inference samples size boosts the likelihood of generating high-quality data. Data synthesis techniques have also been applied to LMMs, improving general vision tasks like visual question answering (Zhou et al., 2024a; Deng et al., 2024; Gao et al., 2023; Zhou et al., 2024b) and enhancing instruction-following capabilities (Luo et al., 2024). Our work extends this line of research by focusing on domain-specific visual classification to enable effective visual assistance in professional tasks. In contrast to existing methods, we address the unique challenges of data synthesis in specialized domains, thereby extending these techniques to support expert-driven applications.